Nine dogs are pulling a sled in a race. The combined mass of the sled and the driver is 220 kg.
The dogs start from rest and reach a speed of 5.0km/h in 8.0 seconds.

What is the average force applied by each dog?

What is the frictional force on the sled if each dog is pulling with a maximum force of 51 N.
1
Expert's answer
2019-12-24T14:28:32-0500

The average force applied by each dog is


"F=\frac{ma}{N}=\frac{mv}{Nt}=\frac{(220)\frac{5}{3.6}}{9(8)}=4.2\ N"

The frictional force on the sled is

"F_f=9(51-4.2)=420\ N"